OpenStudy (anonymous):

It took Malcom 3.5 hours to drive from city A to city B. On his way back to city A, he increased his speed by 20 km per hour and it took him 3 hours. a) Find the average speed for the whole journey.

OpenStudy (anonymous):

Average speed is \[ \frac{\Delta positon}{\Delta time} = \frac{distance\ traveled}{amount\ of\ time\ it\ took} \]

OpenStudy (anonymous):

Which is...

OpenStudy (anonymous):

Hmm, well, come up with some variables...

OpenStudy (anonymous):

Say the traveled distance \(d\) and the going there was speed \(v\).

OpenStudy (anonymous):

\[ 3.5v = d = 3(v+20) \]

OpenStudy (anonymous):

Got it. 129.2 km/hr Right?

OpenStudy (anonymous):

Umm, not sure cause I haven't actually solved it yet.

OpenStudy (anonymous):

Hahahaa :D

OpenStudy (anonymous):

The \(a\) in there is my average velocity.